OpenStudy (anonymous):

A spherical ballon is filled with 4500 pie cubic meters of He gas. If a leak in the ballon causes the gas to escape at the rate of 72pie cubic meters per min find the rate at which (min) the radius of the balloon decreases after 49 min?

OpenStudy (anonymous):

@shubhamsrg @mukushla @hartnn

hartnn (hartnn):

the rate cannot be only in cubic meters....i should be in cubic meters per minute or cubic meters per unit time(like sec,min,hrs)....

OpenStudy (shubhamsrg):

exactly !

OpenStudy (anonymous):

sorry it isin min

hartnn (hartnn):

2nd thing,if the rate does not change with time,whatever be the radius,rate remains same......either 72 is the answer or rate should be the function of time....

hartnn (hartnn):

and as per my guess,rate should decrease with time....

OpenStudy (anonymous):

972 pie

OpenStudy (anonymous):

@magepker728 we have to find the rate of decreasing radius

OpenStudy (shubhamsrg):

well,, initial volume = 4500 after 49 mins, gas escaped = 72 * 49 = 3528 remaining volume = 4500 - 3528 = 472 final radius,r = [ (472))(3)/(4)(pi) ] ^1/3 given => dV/dt = -72 V= 4/3 pi r^3 dV/dt = 4/3 pi (3r^2) dr/dt plug in the values and solve for dr/dt..

OpenStudy (anonymous):